OpenStudy (anonymous):

how do you simplify (7k) ^2

OpenStudy (anonymous):

49k^2

OpenStudy (anonymous):

You distribute the exponent to 7 and k so (7k)^2 = (7^2)(k^2) =49k^2.

OpenStudy (luffingsails):

Again, the question was how to do it. Giving the right answer provides nothing. Nice job, december.